# Sortability of Time Series Data

### v2

@@ Results
We observe very high varsortability values for all datasets; almost all datasets have values over $0.9$ with the wt_pc_validate_v1 dataset beeing the only exception at $0.78$ as shown in Table 3. All $R^{2}$-sortability values are below $0.25$. Most of them are between $0.2$ and $0.32$. The lt_camera_validate_v1, lt_malus_v1, lt_test_v1, wt_bernoulli_v1 and wt_validate_v1 have values very close to 00. The values for each individual experiment contained in one dataset can be found in Appendix C.
PCMCI+ varsortnregress varsortnregressrev. $R^{2}$-sortnregress random
lt_camera_validate_v1 0.08 0.37 0.05 0.05 0.25
lt_malus_v1 0.20 0.24 0.00 0.25 0.19
Table 4: Results for different benchmark algorithms on two Causal Chamber datasets
6 Discussion & Conclusion
In general, both varsortability and $R^{2}$-sortability are present in both simulated and real datasets for benchmarking causal discovery algorithms. In line with Reisach et al. [2021], we observe that DYNOTEARS, which is a NOTEARS based algorithm, seems to perform better when varsortability is higher.
Furthermore, we observe that the data used in the NEURIPS competitionRunge et al. [2019] are highly varsortable, especially the realistic datasets, by our defined time series varsortability metric. The low varsortability in the simulated var models could potentially be due to the fact that these datasets do not have contemporaneous dependencies, and as we showed earlier, contemporaneous dependencies seem to be the driver for high varsortability in simulated time series data. This is hardly surprising given that a team exploiting this effect won the competition[Weichwald et al., 2020]. As for $R^{2}$-sortability, we see that high values can lead to better performance of simple benchmark algorithms. However, $R^{2}$-sortability values observed on the NEURIPS and ER Datasets seem to be too low to be exploited by the sorting algorithm.
@@ Conclusion
In conclusion, our paper represents an empirical extension of the work of [Reisach et al., 2021, 2023]. We demonstrate that high var-sortability occurs in SVAR-simulated time series data, resulting in enhanced performance of score-based causal discovery algorithms assuming equal noise variance. Consequently, we advise caution when assuming equal noise variance for time series causal discovery algorithms. Furthermore, it may be advisable to examine simulated data for high varsortability before using them as benchmark data, as was done in the 2019 NEURIPS competition[Runge et al., 2019]. Finally, we observe high and low varsortability, as well as $R^{2}$-sortability, in two different types of real-life datasets: the Causal Chamber data is generated in a controlled environment while the river flow dataset is measured in an uncontrolled environment. This indicates that var- and $R^{2}$-sortability in auto-correlated time series data is not solely a phenomenon observed in simulated data. For this reason, we believe that marginal variances may contain relevant causal information and exploiting variance or inverse variance sorting may be justified in some situations, if one can combine it with physical reasons for one or the other (even though these physical reasons might already eliminate the need for causal discovery in the first place). Furthermore, the observed $R^{2}$-sortability scores indicate that the assumption of equal noise variance is a tricky one as well. At the very least the relative fraction of unexplained variance may change throughout the graph and unequal noise variances are one possible reason for this.
Acknowledgements.
J.W. received funding from the European Research Council (ERC) Starting Grant CausalEarth under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 948112).

### v3

@@ Results
We observe very high varsortability values for all datasets; almost all datasets have values over $0.9$ with the wt_pc_validate_v1 dataset beeing the only exception at $0.78$ as shown in Table 3. All $R^{2}$-sortability values are below $0.25$. Most of them are between $0.2$ and $0.32$. The lt_camera_validate_v1, lt_malus_v1, lt_test_v1, wt_bernoulli_v1 and wt_validate_v1 have values very close to 0. The values for each individual experiment contained in one dataset can be found in Appendix C. As shown in Table Table 3 varsortnregress outperforms other causal discovery algorithms on 9 out of 10 of the selected Causal Chamber datasets. We only report F1-scores for 10 of the data sets since for For wt_pc_validate_v1 there are not enough samples to select the hyperparameter for sortnregress and the other 6 data sets have over $50,000$ samples resulting in very long execution times.
6 Discussion and Conclusion
In general, both varsortability and $R^{2}$-sortability are present in both simulated and real datasets for benchmarking causal discovery algorithms. In line with Reisach et al. (2021), we observe that DYNOTEARS, which is a NOTEARS based algorithm, seems to perform better when varsortability is higher, which is in line with our hypothesis in section 3.
Furthermore, we observe that the data used in the NEURIPS competition(Runge et al., 2019) are highly varsortable, especially the realistic datasets, by our defined time series varsortability metric. This is also reflected by the good performance of sortnregress, which should be too simple to perform as good as the more sophisticated PCMCI+ algorithm.
The low varsortability in the simulated var models could potentially be due to the fact that these datasets do not have contemporaneous dependencies, and as we showed earlier, contemporaneous dependencies seem to be the driver for high varsortability in simulated time series data. This is hardly surprising given that a team exploiting this effect won the competition(Weichwald et al., 2020). As for $R^{2}$-sortability, we see that high values can lead to better performance of simple benchmark algorithms. However, $R^{2}$-sortability values observed on the NEURIPS and ER Datasets seem to be too low to be exploited by the sorting algorithm.
@@ Conclusion
In conclusion, our paper represents an empirical extension of the work of Reisach et al. (2021; 2023). We demonstrate that high var-sortability occurs in SVAR-simulated time series data, resulting in enhanced performance of continuous score-based causal discovery algorithms assuming equal noise variance. Moreover, in some settings our simple benchmark algorithms outperformed or were en par with more sophisticated algorithms in the presence of high sortability. Consequently, we advise caution when assuming equal noise variance for time series causal discovery algorithms. Furthermore, it may be advisable to examine simulated data for high varsortability before using them as benchmark data, as was done in the 2019 NEURIPS competition(Runge et al., 2019). Finally, we observe high and low varsortability, as well as $R^{2}$-sortability, in two different types of real-life datasets: the Causal Chamber data is generated in a controlled environment while the river flow dataset is measured in an uncontrolled environment. This indicates that var- and $R^{2}$-sortability in auto-correlated time series data is not solely a phenomenon observed in simulated data. For this reason, we believe that marginal variances may contain relevant causal information and exploiting variance or inverse variance sorting may be justified in some situations, if one can combine it with physical reasons for one or the other (even though these physical reasons might already eliminate the need for causal discovery in the first place). Furthermore, the observed $R^{2}$-sortability scores indicate that the assumption of equal noise variance is equally tricky as the relative fraction of unexplained variance may change throughout the graph and unequal noise variances are one possible reason for this.
Acknowledgments
JW was supported by the European Research Council (ERC) Starting Grant CausalEarth under the European Union’s Horizon 2020 research and innovation program (Grant AgreementNo. 948112, PI Prof. Jakob Runge) as well as by the European Regional Development Fund (ERDF) and the German Federal State of Saarland as part of the project (To)CERTAIN. During the initial phase of this project, JW was employed at the Technical University of Berlin and affiliated with the DLR Institute of Data Science as a guest researcher.